# Towards Multimodal Human Intention Understanding Debiasing via Subject-Deconfounding

### v1

@@ 4.2 Model Zoo
Graph-MFN [60] 29.6 75.4 76.6
TFN [59] 32.1 73.9 73.4
LMF [27] 32.8 76.4 75.7
MFM [43] 36.2 78.1 78.1
RAVEN [44] 33.2 78.0 76.6
MCTN [34] 35.6 79.3 79.1
TCSP [45] - 80.9 81.0
PMR [28] 40.6 83.6 83.4
FDMER [48] 42.1 84.2 83.9
MulT [42] 39.5 82.6 82.5
MulT + SuCI 40.7${}^{{\color[rgb]{1,0,0}+1.2}}$ 83.4${}^{{\color[rgb]{1,0,0}+0.8}}$ 83.4${}^{{\color[rgb]{1,0,0}+0.9}}$
MISA [9] 40.2 81.8 81.9
MISA + SuCI ${}^{{\color[rgb]{1,0,0}+1.4}}$ 83.3${}^{{\color[rgb]{1,0,0}+1.5}}$ 83.1${}^{{\color[rgb]{1,0,0}+1.2}}$
Self-MM [58] 41.6 83.9 84.1
Self-MM + SuCI 42.0${}^{{\color[rgb]{1,0,0}+0.4}}$ 84.3${}^{{\color[rgb]{1,0,0}+0.4}}$ 84.6${}^{{\color[rgb]{1,0,0}+0.5}}$
MMIM [12] 41.8 84.1 84.1
MMIM + SuCI 42.4${}^{{\color[rgb]{1,0,0}+0.6}}$ 84.9${}^{{\color[rgb]{1,0,0}+0.8}}$ 84.8${}^{{\color[rgb]{1,0,0}+0.7}}$
DMD [21] 41.0 83.3 83.2
DMD + SuCI 42.2${}^{{\color[rgb]{1,0,0}+1.2}}$ 84.6${}^{{\color[rgb]{1,0,0}+1.3}}$ 84.5${}^{{\color[rgb]{1,0,0}+1.3}}$
4.3 Implementation Details
All models follow consistent feature extraction procedures for fair comparisons. The text feature extractor is instantiated by pre-trained Glove word embedding tool [33] to obtain 300-dimensional linguistic vectors. For MOSI & MOSEI, we use the library Facet [16] to extract an ensemble of visual features, including 35 facial action units to reflect facial gesture changes. Meanwhile, Openface [2] is utilized on UR_FUNNY to extract 75-dimensional features related to facial behaviors and expressions. The audio feature extraction is executed utilizing the software COVAREP [8] to obtain diverse acoustic attributes, where the dimensions on MOSI & MOSEI and UR_FUNNY are 74 and 81, respectively. The word-aligned data points are employed across all benchmarks. In the SuCI implementation, the hidden dimensions $d_{h}$ and $d_{n}$ are set to 64 and 128, respectively. The size $d_{s}$ of each subject confounder is 325, 409, and 456 on MOSI, MOSEI, and UR_FUNNY, respectively. We implement the selected methods and SuCI on NVIDIA Tesla A800 GPUs utilizing the PyTorch toolbox [29], where other training settings are aligned to their original protocols, including the optimizer, additional loss, batch size, learning rate, etc.
4.4 Comparison with State-of-the-art Methods
@@
Graph-MFN [60] 45.0 76.9 77.0
RAVEN [44] 50.0 79.1 79.5
MCTN [34] 49.6 79.8 80.6
TCSP [45] - 82.8 82.7
PMR [28] 52.5 83.3 82.6
FDMER [48] 53.8 83.9 83.8
MulT [42] 51.2 82.1 81.9
MulT + SuCI 52.4${}^{{\color[rgb]{1,0,0}+1.2}}$ 83.2+1.${}^{{\color[rgb]{1,0,0}+1.1}}$ 82.9${}^{{\color[rgb]{1,0,0}+1.0}}$
MISA [9] 51.3 82.3 82.3
MISA + SuCI 52.6${}^{{\color[rgb]{1,0,0}+1.3}}$ 83.5${}^{{\color[rgb]{1,0,0}+1.2}}$ 83.2${}^{{\color[rgb]{1,0,0}+0.9}}$
Self-MM [58] 52.9 83.9 83.8
Self-MM + SuCI 53.6${}^{{\color[rgb]{1,0,0}+0.7}}$ 84.2${}^{{\color[rgb]{1,0,0}+0.3}}$ 84.2${}^{{\color[rgb]{1,0,0}+0.4}}$
MMIM [12] 53.7 84.4 84.3
MMIM + SuCI 54.4${}^{{\color[rgb]{1,0,0}+0.7}}$ 85.5+1.${}^{{\color[rgb]{1,0,0}+1.1}}$ 85.2${}^{{\color[rgb]{1,0,0}+0.9}}$
DMD [21] 53.5 84.1 84.0
DMD + SuCI 54.6+1.${}^{{\color[rgb]{1,0,0}+1.1}}$ 85.8${}^{{\color[rgb]{1,0,0}+1.7}}$ 85.7${}^{{\color[rgb]{1,0,0}+1.7}}$
We compare the SuCI-based models with extensive SOTA methods, including EF-LSTM, LF-LSTM, C-MFN [13], Graph-MFN [60], TFN [59], MFM [43], RAVEN [44], MCTN [34], TCSP [45], PMR [28], and FDMER [48]. The core observations are shown below.
Results on MOSI Benchmark. (i) From Table 1, SuCI consistently improves the performance of the selected methods on all metrics. Concretely, the overall gains of $Acc_{7}$, $Acc_{2}$, and $F1$ scores among the five models increased by 4.8%, 4.8%, and 4.6%, respectively. These improvements confirm that our module can break through the performance bottlenecks of most baselines in a model-agnostic manner. (ii) SuCI can bring more favorable gains for decoupled learning-based efforts. For instance, MISA and DMD achieve salient relative improvements among different metrics of 1.8$\sim$3.5and 1.6$\sim$2.9%, respectively. A reasonable explanation is that the decoupling pattern diffuses the spurious semantics caused by subject confounders in multiple feature subspaces. In this case, SuCI’s de-confounding ability is more effective.
@@
Methods Context Punchline $Acc_{2}\uparrow$ (%)
C-MFN [13] 58.45
C-MFN [13] 64.47
TFN [59] 64.71
LMF [27] 65.16
C-MFN [13] 65.23
FDMER [48] 70.55
MulT [42] 66.65
MulT + SuCI 67.88${}^{{\color[rgb]{1,0,0}+1.23}}$
MISA [9] 67.34
MISA + SuCI 68.96${}^{{\color[rgb]{1,0,0}+1.62}}$
Self-MM [58] 68.77
Self-MM + SuCI 69.72${}^{{\color[rgb]{1,0,0}+0.95}}$
MMIM [12] 69.53
MMIM + SuCI 70.92${}^{{\color[rgb]{1,0,0}+1.39}}$
DMD [21] 68.70
DMD + SuCI 70.84${}^{{\color[rgb]{1,0,0}+2.14}}$
(iii) We also find that the SuCI-based Self-MM is a slight increment. The potential reason is that the unimodal prediction branches of the vanilla model trigger accidental leakage of backdoor effects [30], thus weakening our causal intervention. (iv) Compared to recent PMR and FDMER with complex network stacking and massive parameters [28, 48], MMIM achieves the best results by equipping our efficient SuCI.
Results on MOSEI Benchmark. Table 2 provides performance comparison results on MOSEI. (i) The SuCI-based models outperform the vanilla methods by large margins on all evaluation metrics. For example, SuCI helps the latest DMD (CVPR 2023) to achieve new SOTA performance with considerable absolute gains of 1.1%, 1.7%, and 1.7% in $Acc_{7}$, $Acc_{2}$, and $F1$ scores, respectively. These observations demonstrate the broad applicability and usefulness of our module in removing the subject-related prediction bias. (ii) The improvements provided by SuCI on MOSEI are more significant than those on MOSI. The plausible deduction is that MOSEI contains richer subjects and their highly idiosyncratic multimodal utterances in various scenarios. Thus, SuCI can more accurately remove spurious correlations caused by appropriately extracted subject confounders and offer sufficient gains.
Table 4: Cross-dataset evaluation results of SuCI-based models trained on the MOSI training set and tested on the MOSEI testing set.
Table 5: Cross-dataset evaluation results of SuCI-based models trained on the MOSEI training set and tested on the MOSI testing set.
@@
MulT [42] 46.9 77.6 77.4
MulT + SuCI 48.5${}^{{\color[rgb]{1,0,0}+1.6}}$ 80.2${}^{{\color[rgb]{1,0,0}+2.6}}$ 80.3${}^{{\color[rgb]{1,0,0}+2.9}}$
MISA [9] 46.6 77.2 77.1
MISA + SuCI 48.3${}^{{\color[rgb]{1,0,0}+1.7}}$ 78.9${}^{{\color[rgb]{1,0,0}+1.7}}$ 79.4${}^{{\color[rgb]{1,0,0}+2.3}}$
Self-MM [58] 47.7 78.5 78.3
Self-MM + SuCI 48.2${}^{{\color[rgb]{1,0,0}+0.5}}$ 79.5${}^{{\color[rgb]{1,0,0}+1.0}}$ 79.6${}^{{\color[rgb]{1,0,0}+1.3}}$
MMIM [12] 49.3 79.6 79.3
MMIM + SuCI 51.8${}^{{\color[rgb]{1,0,0}+2.5}}$ 82.5${}^{{\color[rgb]{1,0,0}+2.9}}$ 82.4${}^{{\color[rgb]{1,0,0}+3.1}}$
DMD [21] 48.9 80.8 80.7
DMD + SuCI 51.6${}^{{\color[rgb]{1,0,0}+2.7}}$ 82.4${}^{{\color[rgb]{1,0,0}+1.6}}$ 82.4${}^{{\color[rgb]{1,0,0}+1.7}}$
@@
MulT [42] 37.4 80.2 79.9
MulT + SuCI 38.7${}^{{\color[rgb]{1,0,0}+1.3}}$ 81.7${}^{{\color[rgb]{1,0,0}+1.5}}$ 81.5${}^{{\color[rgb]{1,0,0}+1.6}}$
MISA [9] 37.8 80.5 80.7
MISA + SuCI 39.0${}^{{\color[rgb]{1,0,0}+1.2}}$ 82.2${}^{{\color[rgb]{1,0,0}+1.7}}$ 82.3${}^{{\color[rgb]{1,0,0}+1.6}}$
Self-MM [58] 38.9 82.0 82.1
Self-MM + SuCI 39.5${}^{{\color[rgb]{1,0,0}+0.6}}$ 82.7${}^{{\color[rgb]{1,0,0}+0.7}}$ 82.4${}^{{\color[rgb]{1,0,0}+0.3}}$
MMIM [12] 39.3 82.5 82.5
MMIM + SuCI 41.1${}^{{\color[rgb]{1,0,0}+1.8}}$ 83.4${}^{{\color[rgb]{1,0,0}+0.9}}$ 83.3${}^{{\color[rgb]{1,0,0}+0.8}}$
DMD [21] 38.6 81.6 81.4
DMD + SuCI 40.6${}^{{\color[rgb]{1,0,0}+2.0}}$ 83.0${}^{{\color[rgb]{1,0,0}+1.4}}$ 82.9${}^{{\color[rgb]{1,0,0}+1.5}}$
Table 5: Cross-dataset evaluation results of SuCI-based models trained on the MOSEI training set and tested on the MOSI testing set.
Results on UR_FUNNY Benchmark. As shown in Table 3, we mainly show the detection results using the target punchline for fair comparisons with other works. (i) Intuitively, the SuCI-based models achieve consistent performance increases and accomplish competitive and better results than previous methods. These substantial improvements imply the necessity of performing subject de-confounded training in detecting human humor. (ii) In particular, MMIM and DMD equipped with SuCI yield absolute gains of 1.39% and 2.14%, achieving new SOTAs with the $Acc_{2}$ of 70.92% and 70.84%, respectively.
4.5 Cross-dataset Evaluation
@@
4.6 Ablation Studies
In Table 6, we perform systematic ablation studies to evaluate the impact of all components in SuCI. We report results for the $Acc_{2}$ metrics due to similar trends for the other metrics. The core observations are as follows.
@@ 4.6 Ablation Studies
Table 6: Ablation study results on the three benchmarks. “w/” and “w/o” are short for the with and without, respectively.
@@
MISA [9] DMD [21] MISA [9] DMD [21] MISA [9] DMD [21]
@@
Importance of Different Modalities. (i) When the text, visual, and audio modalities from the subjects are removed separately in de-confounded training, the improvements in SuCI for the baselines show significant deterioration. These decreased results confirm that subject-specific spurious characteristics are transmitted in multimodal utterances and that considering the complete modalities is necessary. (ii) The reason for the pronounced effect of the text modality may be the adverse statistical shortcuts caused by the highly uneven distribution of textual words across the samples of distinct subjects [39], which is a confounder inducer. Also, the impacts of visual and audio modalities are equally essential.
Rationality of Confounder Dictionary. We provide two candidates of the same size as the default confounder dictionary $\bm{Z}$ to evaluate the rationality of our confounder construction. These two alternative dictionaries are obtained by random initialization and unsupervised K-Means++ clustering [1]. (i) The random $\bm{Z}$ would significantly impair the performance and even underperform the vanilla methods, justifying the proposed subject prototypes as confounders. (ii) The performance gains of the clustered $\bm{Z}$ are sub-optimal due to the lack of subject-specific information supervision, leading to the indistinguishability and coupling of confounding effects across different subjects.
@@
Figure 4: (a) Ablation study results for the number of subject confounders on the UR_FUNNY benchmark. Similar trends are observed on other benchmarks. (b) Quantitative results (i.e., binary or seven classifications) of vanilla and SuCI-based DMD.
Impact of Subject Confounder Number. Finally, we test the impact of different numbers of subject confounders on SuCI performance in Fig. 4(a). The SuCI-based models all exhibit overall rising gain trends as the training subjects increase (i.e., the number of used human subjects from 100 to 1,000). These findings suggest that sufficient stratified confounders facilitate better backdoor adjustment implementation and accurate average causal effect estimation.
4.7 Qualitative Evaluation
To show the difference between the model approximate $P(\bm{Y}|\bm{X})$ and $P(\bm{Y}|do(\bm{X}))$, we randomly select 20 samples in each sentiment category on MOSEI for visualization. As Fig. 4(b) shows, the sample distributions from the different categories of the vanilla baseline prediction usually appear confounded and fragmented. In comparison, the causality-based SuCI improves the predicted results so that the samples in the same category are more compact and different categories are well separated. The observation suggests that our causal intervention promotes discriminative predictions by mitigating the subject prediction bias.
5 Conclusion
This paper is the first to reveal the long-neglected subject variation problem in MIU tasks and identify subjects as essentially harmful confounders from a causal perspective. Thus, we present a subject causal intervention module (SuCI) to remove the prediction bias caused by the subject-specific spurious correlations. Extensive experiments show the broad applicability of SuCI in diverse methods.

### v3

@@ Introduction
Graph-MFN 29.6 75.4 76.6
TFN 32.1 73.9 73.4
LMF 32.8 76.4 75.7
MFM 36.2 78.1 78.1
RAVEN 33.2 78.0 76.6
MCTN 35.6 79.3 79.1
TCSP - 80.9 81.0
PMR 40.6 83.6 83.4
FDMER 42.1 84.2 83.9
MulT 39.5 82.6 82.5
MulT + SuCI 40.7+1.2 83.4+0.8 83.4+0.9
MISA 40.2 81.8 81.9
MISA + SuCI 41.6+1.4 83.3+1.5 83.1+1.2
Self-MM 41.6 83.9 84.1
Self-MM + SuCI 42.0+0.4 84.3+0.4 84.6+0.5
MMIM 41.8 84.1 84.1
MMIM + SuCI 42.4+0.6 84.9+0.8 84.8+0.7
DMD 41.0 83.3 83.2
DMD + SuCI 42.2+1.2 84.6+1.3 84.5+1.3
Benchmarks and Model Zoo
Benchmarks. Extensive experiments are conducted on three mainstream MLU benchmarks. Concretely, MOSI (Zadeh et al. 2016) is a multimodal human sentiment analysis dataset consisting of 2,199 video segments. The standard data partitioning is 1,284 samples for training, 284 samples for validation, and 686 samples for testing. These samples contain a total of 89 distinct subjects from video blogs. Each sample is manually annotated with a sentiment score ranging from -3 to 3. MOSEI (Zadeh and Pu 2018) is a large-scale human sentiment and emotion recognition benchmark containing 22,856 video clips from 1,000 different subjects and 250 diverse topics. Among these samples, 16,326, 1,871, and 4,659 samples are used as training, validation and testing sets. The sample annotation protocols used for sentiment analysis are consistent with MOSI. UR_FUNNY (Hasan et al. 2019) is a multimodal human humor detection dataset that contains 16,514 video clips from 1,741 subjects collected by the TED portal. There are 10,598, 2,626, and 3,290 samples in the training, validation, and testing sets. Each sample provides the target punchlines and associated contexts from multimodal utterances to support the detection of subject humor/non-humor in the binary labels. Following widely adopted implementations (Yang et al. 2022a; Li, Wang, and Cui 2023), we use the 7-class accuracy ($Acc_{7}$), binary accuracy ($Acc_{2}$), and $F1$ score to evaluate the results on MOSI and MOSEI. The standard binary accuracy ($Acc_{2}$) is utilized for evaluation on UR_FUNNY.
@@
Graph-MFN 45.0 76.9 77.0
RAVEN 50.0 79.1 79.5
MCTN 49.6 79.8 80.6
TCSP - 82.8 82.7
PMR 52.5 83.3 82.6
FDMER 53.8 83.9 83.8
MulT 51.2 82.1 81.9
MulT + SuCI 52.4+1.2 83.2+1.1 82.9+1.0
MISA 51.3 82.3 82.3
MISA + SuCI 52.6+1.3 83.5+1.2 83.2+0.9
Self-MM 52.9 83.9 83.8
Self-MM + SuCI 53.6+0.7 84.2+0.3 84.2+0.4
MMIM 53.7 84.4 84.3
MMIM + SuCI 54.4+0.7 85.5+1.1 85.2+0.9
DMD 53.5 84.1 84.0
DMD + SuCI 54.6+1.1 85.8+1.7 85.7+1.7
Model Zoo. Considering the applicability, we choose five representative models with different network structures to verify the proposed plug-in SuCI. A brief overview is as follows. MulT (Tsai et al. 2019) constructs multimodal transformers to learn element dependencies between pairwise modalities. MISA (Devamanyu, Roger, and Soujanya 2020) captures modality-invariant and modality-specific diverse representations based on feature disentanglement. Self-MM (Yu et al. 2021) utilizes a self-supervised paradigm to learn additional emotion semantics from unimodal label generation. MMIM (Han, Chen, and Poria 2021) proposes a hierarchical mutual information maximization to alleviate the loss of valuable task-related clues. DMD (Li, Wang, and Cui 2023) designs cross-modal knowledge distillations to bridge the semantic gap among modalities.
Implementation Details
All models follow consistent feature extraction procedures for fair comparisons. The text feature extractor is instantiated by pre-trained Glove word embedding tool (Pennington, Socher, and Manning 2014) to obtain 300-dimensional linguistic vectors. For MOSI & MOSEI, we use the library Facet (iMotions 2017) to extract an ensemble of visual features, including 35 facial action units to reflect facial gesture changes. Meanwhile, Openface (Baltrušaitis, Robinson, and Morency 2016) is utilized on UR_FUNNY to extract 75-dimensional features related to facial behaviors and expressions. The audio feature extraction is executed utilizing the software COVAREP (Degottex et al. 2014) to obtain diverse acoustic attributes, where the dimensions on MOSI & MOSEI and UR_FUNNY are 74 and 81, respectively. The word-aligned data points are employed across all benchmarks. In the SuCI implementation, the hidden dimensions $d_{h}$ and $d_{n}$ are set to 64 and 128, respectively. The size $d_{s}$ of each subject confounder is 325, 409, and 456 on MOSI, MOSEI, and UR_FUNNY, respectively. We implement the selected methods and SuCI on NVIDIA Tesla A800 GPUs utilizing the PyTorch toolbox, where other training settings are aligned to their original protocols.
Methods Context Punchline $Acc_{2}\uparrow$ (%)
C-MFN 58.45
C-MFN 64.47
TFN 64.71
LMF 65.16
C-MFN 65.23
FDMER 70.55
MulT 66.65
MulT + SuCI 67.88+1.23
MISA 67.34
MISA + SuCI 68.96+1.62
Self-MM 68.77
Self-MM + SuCI 69.72+0.95
MMIM 69.53
MMIM + SuCI 70.92+1.39
DMD 68.70
DMD + SuCI 70.84+2.14
@@
Comparison with State-of-the-art Methods
We compare the SuCI-based models with extensive SOTA methods, including EF-LSTM, LF-LSTM, C-MFN (Hasan et al. 2019), Graph-MFN (Zadeh and Pu 2018), TFN (Zadeh et al. 2017), MFM (Tsai et al. 2018), RAVEN (Wang et al. 2019), MCTN (Pham et al. 2019), TCSP (Wu et al. 2021), PMR (Lv et al. 2021), and FDMER (Yang et al. 2022a).
Results on MOSI Benchmark. (i) From Table 1, SuCI consistently improves the performance of the selected methods on all metrics. Concretely, the overall gains of $Acc_{7}$, $Acc_{2}$, and $F1$ scores among the five models increased by 4.8%, 4.8%, and 4.6%, respectively. These improvements confirm that our module can break through the performance bottlenecks of most baselines in a model-agnostic manner. (ii) SuCI can bring more favorable gains for decoupled learning-based efforts. For instance, MISA and DMD achieve salient relative improvements among different metrics of 1.8$\sim$3.5and 1.6$\sim$2.9%, respectively. A reasonable explanation is that the decoupling pattern diffuses the spurious semantics caused by subject confounders in multiple feature subspaces. In this case, SuCI’s de-confounding ability is more effective. (iii) Compared to recent PMR and FDMER with complex network stacking and massive parameters (Lv et al. 2021; Yang et al. 2022a), MMIM achieves the best results by equipping our SuCI.
@@
MulT 46.9 77.6 77.4
MulT + SuCI 48.5+1.6 80.2+2.6 80.3+2.9
MISA 46.6 77.2 77.1
MISA + SuCI 48.3+1.7 78.9+1.7 79.4+2.3
Self-MM 47.7 78.5 78.3
Self-MM + SuCI 48.2+0.5 79.5+1.0 79.6+1.3
MMIM 49.3 79.6 79.3
MMIM + SuCI 51.8+2.5 82.5+2.9 82.4+3.1
DMD 48.9 80.8 80.7
DMD + SuCI 51.6+2.7 82.4+1.6 82.4+1.7
Table 4: Cross-dataset evaluation of models trained on the MOSI training set and tested on the MOSEI testing set.
@@
MulT 37.4 80.2 79.9
MulT + SuCI 38.7+1.3 81.7+1.5 81.5+1.6
MISA 37.8 80.5 80.7
MISA + SuCI 39.0+1.2 82.2+1.7 82.3+1.6
Self-MM 38.9 82.0 82.1
Self-MM + SuCI 39.5+0.6 82.7+0.7 82.4+0.3
MMIM 39.3 82.5 82.5
MMIM + SuCI 41.1+1.8 83.4+0.9 83.3+0.8
DMD 38.6 81.6 81.4
DMD + SuCI 40.6+2.0 83.0+1.4 82.9+1.5
Table 5: Cross-dataset evaluation of models trained on the MOSEI training set and tested on the MOSI testing set.
Results on MOSEI Benchmark. Table 2 provides performance comparison results on MOSEI. (i) The SuCI-based models outperform the vanilla methods by large margins on all metrics. For example, SuCI helps the latest DMD to achieve new SOTA performance with considerable absolute gains of 1.1%, 1.7%, and 1.7% in $Acc_{7}$, $Acc_{2}$, and $F1$ scores, respectively. These observations show the broad applicability and usefulness of our module in removing the subject-related prediction bias. (ii) The improvements provided by SuCI on MOSEI are more significant than those on MOSI. The plausible deduction is that MOSEI contains richer subjects and their highly idiosyncratic multimodal utterances in various scenarios. Thus, SuCI can more accurately remove spurious correlations caused by appropriately extracted subject confounders and offer sufficient gains.
Results on UR_FUNNY Benchmark. From Table 3, we show the detection results using the target punchline for fair comparisons with other works. (i) The SuCI-based models achieve consistent performance increases and accomplish competitive and better results than previous methods. These substantial improvements imply the necessity of performing subject de-confounded training in detecting human humor. (ii) In particular, MMIM and DMD equipped with SuCI yield absolute gains of 1.39% and 2.14%, achieving new SOTAs with the $Acc_{2}$ of 70.92% and 70.84%, respectively.
Cross-dataset Evaluation
@@
Ablation Studies
In Table 6, we perform ablation studies to evaluate the impact of all components in SuCI. We report results for the $Acc_{2}$ metrics due to similar trends for the other metrics.
@@
MISA DMD MISA DMD MISA DMD
@@
Table 6: Ablation studies on the three benchmarks. “w/” and “w/o” mean the with and without.
Importance of Different Modalities. (i) When the text, visual, and audio modalities from the subjects are removed separately in de-confounded training, the improvements in SuCI for the baselines show significant deterioration. These decreased results confirm that subject-specific spurious characteristics are transmitted in multimodal utterances and that considering the complete modalities is necessary. (ii) The reason for the pronounced effect of the text modality may be the adverse statistical shortcuts caused by the highly uneven distribution of textual words across the samples of distinct subjects, which is a confounder inducer.
Figure 4: Ablation study results for the number of subject confounders on the UR_FUNNY benchmark.
Figure 5: Quantitative results (i.e., binary or seven classifications) of vanilla and SuCI-based DMD on the MOSEI.
Rationality of Confounder Dictionary. We provide two candidates of the same size as the default confounder dictionary $\bm{Z}$ to evaluate the rationality of our confounder construction. These two alternative dictionaries are obtained by random initialization and unsupervised K-Means++ clustering (Bahmani et al. 2012). (i) The random $\bm{Z}$ would significantly impair the performance and even underperform the vanilla methods, justifying the proposed subject prototypes as confounders. (ii) The performance gains of the clustered $\bm{Z}$ are sub-optimal due to the lack of subject-specific information supervision, leading to the indistinguishability and coupling of confounding effects across different subjects.
@@
Impact of Subject Confounder Number. Finally, we test the impact of different numbers of subject confounders on SuCI performance in Figure 4. The SuCI-based models all exhibit overall rising gain trends as the training subjects increase. These findings suggest that sufficient stratified confounders facilitate better backdoor adjustment implementation and accurate average causal effect estimation.
Qualitative Evaluation
To show the difference between the model approximate $P(\bm{Y}|\bm{X})$ and $P(\bm{Y}|do(\bm{X}))$, we randomly select several samples in each sentiment category on MOSEI for visualization. As Figure 5 shows, the sample distributions from the different categories of the vanilla baseline prediction usually appear confounded and fragmented. In comparison, the causality-based SuCI improves the predicted results so that the samples in the same category are more compact and different categories are well separated. The observation suggests that our causal intervention promotes discriminative predictions by mitigating the subject prediction bias.
Conclusion
This paper is the first to reveal the long-neglected subject variation problem in MLU tasks and identify subjects as essentially harmful confounders from a causal perspective. Thus, we present a subject causal intervention module (SuCI) to remove the prediction bias caused by the subject-specific spurious correlations. Extensive experiments show the broad applicability of SuCI in diverse methods.
Acknowledgments
This work is supported in part by the National Key R&D Program of China (2021ZD0113503).